# Unlocking Exploration in RLVR: Uncertainty-aware Advantage Shaping for Deeper Reasoning

### v1

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) has shown significant promise for enhancing the reasoning capabilities of large language models (LLMs). However, prevailing algorithms like GRPO broadcast a uniform advantage signal across all tokens in a sequence. This coarse-grained approach overlooks the pivotal role of uncertain, high-stakes decisions during reasoning, leading to inefficient exploration and the well-documented problem of entropy collapse. To address this, we introduce UnCertainty-aware Advantage Shaping (UCAS), a model-free method that refines credit assignment by leveraging the model’s internal uncertainty signals. UCAS operates in two stages: it first modulates the response-level advantage using the model’s overall self-confidence, and then applies a token-level penalty based on raw logit certainty. This dual mechanism encourages exploration of high-uncertainty paths that yield correct answers while penalizing overconfident yet erroneous reasoning, effectively balancing the exploration-exploitation trade-off. Extensive experiments on five mathematical reasoning benchmarks show that UCAS significantly outperforms strong RLVR baselines across multiple model scales, including 1.5B and 7B. Our analysis confirms that UCAS not only achieves higher rewards but also promotes greater reasoning diversity and successfully mitigates entropy collapse. The code and models of UCAS are available at https://github.com/xvolcano02/UCAS.
@@ 1 Introduction
However, the success of RLVR reveals a critical underlying tension: the trade-off between precision and diversity. While methods like GRPO excel at increasing the probability of generating correct answers, they often do so at the cost of exploration. Due to the absence of a critic model, the learning signal in GRPO, which applies a single uniform advantage across all tokens, provides an indiscriminate and overly coarse form of credit assignment. It rewards all steps of a correct path equally and penalizes all steps of an incorrect one, failing to distinguish crucial reasoning leaps from trivial ones. This coarse-grained feedback drives the policy to converge prematurely on a small set of ”safe” high-reward trajectories. A common side effect is entropy collapse (Cui et al., 2025b), where the output distribution contracts, reducing solution diversity and impairing performance on complex problems that demand novel reasoning strategies.
@@
To address the above challenges, we propose an UnCertainty-aware Advantage Shaping (UCAS), a model-free method that refines credit assignment in RLVR by leveraging the model’s intrinsic uncertainty. UCAS is designed to resolve the precision–diversity dilemma by reshaping the advantage signal at two complementary levels. At the response level, UCAS modulates the sequence-level advantage using the model’s overall self-confidence, amplifying rewards for correct-but-uncertain responses and penalties for incorrect-but-confident ones. At the token level, it further introduces a certainty-based penalty derived directly from raw logits, discouraging local overconfidence while preserving diversity in reasoning. Collectively, these mechanisms promote exploration of uncertain but potentially fruitful reasoning paths, while efficiently suppressing confidently wrong solutions. Extensive experiments on five mathematical reasoning benchmarks demonstrate that UCAS consistently outperforms strong RLVR baselines at both the 1.5B and 7B model scales. Beyond reward improvements, UCAS fosters greater reasoning diversity and substantially mitigates entropy collapse, confirming the effectiveness of uncertainty as a fine-grained learning signal.
@@ 2 Background: Reinforcement Learning with Verifiable Rewards
To overcome these limitations, recent research has introduced RLVR. RLVR converts open-ended outputs into programmatically checkable signals, such as numerical consistency in mathematics, unit-test pass rates in code generation, or formal constraint satisfaction (Su et al., 2025; Wang et al., 2025b), thereby avoiding the noise and cost of preference models. By forming a closed loop of model–environment–verifier, RLVR enables policies to be updated directly from binary or graded correctness signals, improving both sample efficiency and reproducibility in structured reasoning tasks.
@@ 3.1 Uncertainty Signals: From Confidence to Logits
For a high-level assessment of a full reasoning trajectory, we use the model’s self-confidence. As defined in Kang et al. (2025), this is the average KL-divergence between the model’s predictive distribution and a uniform distribution over the vocabulary . We denote this as :
A higher value of signifies higher overall confidence (low uncertainty) in the generated sequence, suggesting the model is following a well-trodden path.
@@ 3.2 UCAS: Two-Stage Advantage Shaping
where and are the mean and standard deviation of confidence scores across the group.
We then compute a modulation weight based on the sign of the original advantage , which directly encodes the correctness of the answer.
@@
Response-level modulation sets a global learning objective for each trajectory, but this modulated advantage, , is still a uniform signal broadcast to all tokens within that sequence. This alone is insufficient to prevent the model from developing localized overconfidence—a key driver of entropy collapse. The second stage therefore introduces a token-specific penalty to directly address this. By penalizing high-certainty choices at each step, we encourage the model to maintain a degree of epistemic humility, which preserves local exploration.
@@
where is a hyperparameter controlling the penalty strength. This composite structure steers the model toward novel correct solutions (via ) while ensuring it traverses reasoning paths with a healthy degree of caution (via the penalty term), thereby mitigating entropy collapse and fostering more robust problem-solving abilities. This final advantage term then replaces the original advantage in the RL objective:
@@ 4.1 Experimental Setup
Models and Baselines.
We employ two variants of the Qwen2.5-Math (Yang et al., 2024) series as our foundation models: Qwen2.5-Math-1.5B and Qwen2.5-Math-7B. First, to quantify the performance improvement introduced by our method, we select the widely used GRPO and DAPO algorithms as comparison baselines. Furthermore, to benchmark against existing reinforcement learning techniques, we also select the following representative methods for comparison:
Implementation Details.
We adopt the VERL framework (Sheng et al., 2024) and train our model using the optimization objective defined in Eq. 10. During training, the model’s maximum context length is set to 4096, with a maximum prompt length of 1024 and a maximum response length of 3072. The learning rate is fixed at , and the training batch size is set to 512. For each prompt, we sample 16 rollouts with a sampling temperature of 1.0. For the DAPO baseline, we use clipping thresholds of and . The KL penalty loss and entropy regularization loss are omitted from the objective function. The hyperparameters for our UCAS method, and , are set to 0.25 and 0.01, respectively. All experiments are conducted on 2 compute nodes, each equipped with 8 NVIDIA A800 80GB GPUs.
@@ 4.2 Main Results
The greedy pass@1 performance comparison between 1.5B and 7B models across five mathematical reasoning benchmarks is presented in Table 1. We can clearly find that the UCAS model achieved the highest performance across all five math reasoning benchmarks on both the 1.5B and 7B parameter scales. Compared with the DAPO baseline, UCAS improves the average accuracy from 41.2 to 47.3 (+6.1) on Qwen2.5-Math-1.5B and from 50.5 to 56.7 (+6.2) on Qwen2.5-Math-7B. Beyond DAPO, UCAS also surpasses strong baselines such as KTAE and Oat-Zero, with pronounced gains on challenging benchmarks including AIME24, AMC, and OlympiadBench. These results highlight the robustness and scalability of uncertainty-aware advantage shaping, demonstrating consistent benefits across model sizes and diverse reasoning tasks.
@@ 4.3.1 Ablation Study
4.3.2 Training Dynamics
@@
4.3.3 Pass@k Evaluation
Prior studies (Wang et al., 2022; Wu et al., 2024) have shown that with a limited number of rollouts, models often struggle to solve certain tasks. In contrast, when the rollout budget is sufficiently large, the probability of sampling effective solutions increases considerably. This observation suggests that pass@k accuracy with a large k provides a more reliable estimate of a model’s potential performance (Yue et al., 2025). Under this evaluation protocol, a problem is considered solved if any of the k sampled reasoning trajectories yield the correct answer.
Figure 4 reports pass@k results on the AIME24 benchmark. The results indicate that UCAS achieves more consistent improvements as k grows. In contrast, Vanilla-GRPO and its enhanced variants show slower growth, consistent with findings from Yue et al. (2025). The stronger performance of UCAS under the pass@k metric highlights its effectiveness, which can be attributed to differences in exploration strategies. Unlike Vanilla-GRPO, which often suffers from exploration stagnation, where the model repeatedly samples low-diversity rollouts, UCAS leverages uncertainty-aware advantage shaping to sustain diverse exploration and escape local optima.
@@ 5.1 RL for LLM Reasoning
Recent advances in reinforcement learning have transformed the training of large language models for reasoning tasks. Process reward models (PRMs) (Lightman et al., 2024) have emerged as a key innovation, providing step-level supervision that improves both efficiency and accuracy compared to outcome-only rewards. Approaches such as PRIME (Cui et al., 2025a) eliminate costly human annotation by deriving implicit process feedback, while OmegaPRM (Luo et al., 2024) leverages Monte Carlo Tree Search (MCTS) to automatically identify reasoning errors. Alongside this, DeepSeek-R1 (Guo et al., 2025) demonstrates that sophisticated reasoning can emerge purely from RL without supervised fine-tuning, enabled by GRPO, which replaces value functions with group-based baselines. These advances redefine alignment and reasoning in LLMs, positioning reinforcement learning with verifiable or process-level rewards as a scalable and principled alternative to preference-model-based RLHF.
@@ 5.2 Reinforcement Learning from Verifiable Rewards
RLVR has emerged as a scalable alternative to preference-based alignment by converting open-ended outputs into checkable signals such as mathematical correctness or unit-test pass rates (Guo et al., 2025; Yue et al., 2025). While early implementations demonstrated strong gains in pass@1 accuracy, subsequent studies revealed a consistent policy entropy collapse: models rapidly concentrate probability mass on a narrow set of high-reward trajectories, diminishing output diversity and limiting exploration (Cui et al., 2025b). Empirical analyses show that RLVR-trained models often underperform base models on pass@ (Shao et al., 2024; Yue et al., 2025), highlighting a precision–diversity trade-off (Wu et al., 2025; Dong et al., 2025).
Algorithmic responses to entropy collapse vary. Entropy or KL penalties provide partial remedies, though their effectiveness depends on the divergence form (Li et al., 2025). More principled adaptations include token-level covariance control, which selectively limits updates on high-confidence tokens to sustain exploration (Cui et al., 2025b). In parallel, hybrid optimization strategies, exemplified by RL-PLUS (Dong et al., 2025), exploit off-policy data to reintroduce diversity, underscoring entropy preservation as a fundamental bottleneck in scaling RLVR.
@@ 6 Conclusion
In this work, we introduced UnCertainty-aware Advantage Shaping (UCAS), a fine-grained advantage estimation framework that leverages internal confidence signals without requiring additional reward models. By jointly modeling uncertainty at both the response and token levels, UCAS reshapes advantages to highlight critical uncertain reasoning steps and suppress overconfident yet erroneous segments. Experimental results on major mathematical reasoning benchmarks show that UCAS achieves substantial performance improvements over GRPO and its enhanced variants. Analysis of the training dynamics further reveals that, as training progresses, UCAS demonstrates steadily increasing rewards, longer reasoning chains, and an entropy trajectory that first declines and then rises, reflecting stronger exploratory capability. These findings indicate that uncertainty-aware advantage shaping offers an effective pathway toward more robust reinforcement learning for large language models.
Appendix A Limitation and Future Work
While our work demonstrates the effectiveness of UCAS in the domain of mathematical reasoning, we acknowledge several limitations that present valuable opportunities for future research.
First, our experiments are exclusively focused on mathematical tasks, which benefit from clear, binary verifiable rewards (i.e., the answer is either correct or incorrect). The direct applicability of UCAS to domains with more nuanced, subjective, or dense reward signals—such as creative writing, summarization, or open-domain dialogue—remains an open question. Adapting the uncertainty-aware shaping mechanism to these softer reward landscapes would be a crucial next step.
Second, our method relies on self-confidence and raw logits as proxies for model uncertainty. While these internal signals are computationally efficient and effective, future work could explore alternative or complementary uncertainty metrics. Techniques such as Monte Carlo dropout, model ensembles, or semantic entropy could potentially capture different facets of model uncertainty and lead to even more refined and robust advantage shaping. Investigating these areas will be essential for understanding the broader generalizability of our approach.

### v2

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) has shown significant promise for enhancing the reasoning capabilities of large language models (LLMs). However, prevailing algorithms like GRPO broadcast a uniform advantage signal across all tokens in a sequence. This coarse-grained approach overlooks the pivotal role of uncertain, high-stakes decisions during reasoning, leading to inefficient exploration and the well-documented problem of entropy collapse. To address this, we introduce UnCertainty-aware Advantage Shaping (UCAS), a model-free method that refines credit assignment by leveraging the model’s internal uncertainty signals. UCAS operates in two stages: it first modulates the response-level advantage using a logit-space self-confidence proxy, and then applies an asymmetric token-level penalty based on raw logit certainty. This dual mechanism encourages exploration of high-uncertainty paths that yield correct answers while penalizing overconfident yet erroneous reasoning, effectively balancing the exploration-exploitation trade-off. Extensive experiments on five mathematical reasoning benchmarks show that UCAS significantly outperforms strong RLVR baselines across multiple model scales, including 1.5B and 7B. Our analysis confirms that UCAS not only achieves higher rewards but also promotes greater reasoning diversity and successfully mitigates entropy collapse. Code is available at https://github.com/xvolcano02/UCAS.
Unlocking Exploration in RLVR: Uncertainty-aware Advantage Shaping for Deeper Reasoning
Can Xie, Ruotong Pan, Xiangyu Wu, Yunfei Zhang, Jiayi Fu, Tingting Gao, Guorui Zhou††thanks: Corresponding author Kuaishou Technology xiecan02@gmail.com {panruotong03, wuxiangyu06, zhangyunfei05, fujiayi, lisize, zhouguorui}@kuaishou.com
@@ 1 Introduction
However, the success of RLVR reveals a critical underlying tension: the trade-off between precision and diversity. While methods like GRPO excel at increasing the probability of generating correct answers, they often do so at the cost of exploration. Due to the absence of a critic model, the learning signal in GRPO, which applies a single uniform advantage across all tokens, provides an indiscriminate and overly coarse form of credit assignment. It rewards all steps of a correct path equally and penalizes all steps of an incorrect one, failing to distinguish crucial reasoning leaps from trivial ones. This coarse-grained feedback drives the policy to converge prematurely on a small set of "safe" high-reward trajectories. A common side effect is entropy collapse (Cui et al., 2025b), where the output distribution contracts, reducing solution diversity and impairing performance on complex problems that demand novel reasoning strategies.
@@
To address the above challenges, we propose an UnCertainty-aware Advantage Shaping (UCAS), a model-free method that refines credit assignment in RLVR by leveraging the model’s intrinsic uncertainty. UCAS is designed to resolve the precision–diversity dilemma by reshaping the advantage signal at two complementary levels. At the response level, UCAS modulates the sequence-level advantage using a logit-space self-confidence proxy, amplifying rewards for correct-but-uncertain responses and penalties for incorrect-but-confident ones. At the token level, it further introduces an asymmetric certainty-based penalty derived directly from raw logits, suppressing overconfident errors while safeguarding positive reasoning trajectories from negative token-level updates. Collectively, these mechanisms promote exploration of uncertain but potentially fruitful reasoning paths, while efficiently suppressing confidently wrong solutions. Extensive experiments on five mathematical reasoning benchmarks demonstrate that UCAS consistently outperforms strong RLVR baselines at both the 1.5B and 7B model scales. Beyond reward improvements, UCAS fosters greater reasoning diversity and substantially mitigates entropy collapse, confirming the effectiveness of uncertainty as a fine-grained learning signal.
@@ 2 Background: Reinforcement Learning with Verifiable Rewards
To overcome these limitations, recent research has introduced RLVR. RLVR converts open-ended outputs into programmatically checkable signals, such as numerical consistency in mathematics, unit-test pass rates in code generation, or formal constraint satisfaction (Su et al., 2025; Wang et al., 2025b; Lin et al., 2026; Li et al., 2026; Jiang et al., 2026), thereby avoiding the noise and cost of preference models. By forming a closed loop of model–environment–verifier, RLVR enables policies to be updated directly from binary or graded correctness signals, improving both sample efficiency and reproducibility in structured reasoning tasks.
@@ 3.1 Uncertainty Signals: From Confidence to Logits
For a high-level assessment of a full reasoning trajectory, we use a logit-space self-confidence score. Let denote the pre-softmax logit of vocabulary item at step for response , and define the token-level confidence as
where . We then define sequence-level confidence as
This formulation avoids probability-space instability when token probabilities approach zero. As we show in Appendix C, is equivalent to , so larger values of indicate higher overall confidence and lower uncertainty. In contrast, when we discuss entropy collapse later in the paper, we use entropy or as a macro-level diagnostic of distributional concentration. The two directions therefore play complementary roles: serves as the training-time modulation signal, whereas characterizes the collapse outcome.
@@ 3.2 UCAS: Two-Stage Advantage Shaping
where and are the mean and standard deviation of confidence scores across the group. This standardization makes the modulation weights comparable within each prompt group and reduces dependence on model-specific raw logit scales.
We then compute a modulation weight based on the sign of the original advantage , which directly encodes the correctness of the answer. Theoretically, we select an exponential form to act as a non-linear filter. This addresses the compressed variance often found in group-normalized scores, where linear rescaling fails to sufficiently distinguish “novel” exploration from “routine” exploitation. More theoretical explanation can be found in Appendix C.
@@
Response-level modulation sets a global learning objective for each trajectory, but this modulated advantage, , is still a uniform signal broadcast to all tokens within that sequence. This alone is insufficient to prevent the model from developing localized overconfidence—a key driver of entropy collapse. The second stage therefore introduces a token-specific penalty to directly address this. The penalty is applied asymmetrically: it strongly suppresses locally overconfident tokens on unfavorable trajectories, while positive trajectories are protected from being flipped into negative token-level updates.
@@
where is a hyperparameter controlling the penalty strength. When , the clamp prevents correct trajectories from receiving negative token-level advantages. When , the penalty further suppresses high-certainty errors, assigning stronger negative updates to confidently wrong transitions. This asymmetric structure steers the model toward novel correct solutions while discouraging localized overconfidence, thereby mitigating entropy collapse and fostering more robust problem-solving abilities. This final advantage term then replaces the original advantage in the RL objective:
@@ 4.1 Experimental Setup
Models and Baselines.
We employ two variants of the Qwen2.5-Math (Yang et al., 2024) series as our foundation models: Qwen2.5-Math-1.5B and Qwen2.5-Math-7B. To quantify the performance improvement introduced by our method, we compare against two widely used RLVR baselines, GRPO and DAPO. In addition, we benchmark against several representative recent methods in math reasoning and RLVR, including Simple-RL-Zoo (Zeng et al., 2025), PRIME-Zero (Cui et al., 2025a), OpenReasonerZero (Hu et al., 2025), Oat-Zero (Liu et al., 2025b), GRPO with Entropy Adv. (Cheng et al., 2025), and KTAE (Sun et al., 2025b). Detailed descriptions of the baselines are provided in Appendix A.1.
@@ 4.2 Main Results
The greedy pass@1 performance comparison between 1.5B and 7B models across five mathematical reasoning benchmarks is presented in Table 1. We can clearly find that the UCAS model achieved the highest performance across all five math reasoning benchmarks on both the 1.5B and 7B parameter scales. Compared with the DAPO baseline, UCAS improves the average accuracy from 41.2 to 47.3 (+6.1) on Qwen2.5-Math-1.5B and from 50.5 to 56.7 (+6.2) on Qwen2.5-Math-7B. Beyond DAPO, UCAS also surpasses strong baselines such as KTAE and Oat-Zero, with pronounced gains on challenging benchmarks including AIME24, AMC, and OlympiadBench. These results highlight the robustness and scalability of uncertainty-aware advantage shaping, demonstrating consistent benefits across model sizes and diverse reasoning tasks. Additional four-seed results in Appendix B.4 show that these gains are stable and not driven by a single favorable run.
@@ 4.3 Analysis
| Response-Level | Token-Level | Math-500 |
| Varying Token-Level Penalty (), Fixed | ||
| 0.2 | 0.005 | 79.6 |
| 0.2 | 0.01 | 80.4 |
| 0.2 | 0.05 | 78.8 |
| 0.2 | 0.1 | 78.4 |
| Varying Response Modulation (), Fixed | ||
| 0.1 | 0.01 | 80.0 |
| 0.2 | 0.01 | 80.4 |
| 0.4 | 0.01 | 78.8 |
Hyperparameter Sensitivity.
We evaluate the robustness of UCAS by varying the response-level modulation and token-level penalty on the Math-500 benchmark (Table 3). This stability is supported by strict normalization throughout the method: group advantages and response-level confidence are z-score normalized, while token logits are min–max normalized to . A moderate penalty () yields optimal performance. Lower values (0.005) fail to sufficiently counteract entropy collapse, while excessive penalization () over-regularizes the policy. This suggests that high artificially flattens the distribution even for necessary, high-certainty steps, hindering coherent reasoning chains. The method exhibits stability within . However, aggressive modulation () causes a performance drop. This implies a "signal dominance" issue: overly strong confidence scaling overshadows the fundamental correctness signal, introducing variance that distracts from the primary objective of mathematical accuracy. Notably, the same setting also transfers effectively across the 1.5B and 7B models in Table 1, indicating limited sensitivity to moderate model-scale variation.
Training Dynamics.
@@
Pass@k Evaluation.
Prior studies (Wang et al., 2022; Wu et al., 2024) have shown that with a limited number of rollouts, models often struggle to solve certain tasks. In contrast, when the rollout budget is sufficiently large, the probability of sampling effective solutions increases considerably. This observation suggests that pass@k accuracy with a large k provides a more reliable estimate of a model’s potential performance (Yue et al., 2025). Under this evaluation protocol, a problem is considered solved if any of the k sampled reasoning trajectories yield the correct answer. Figure 3 reports pass@k results on the AIME24 benchmark. The results indicate that UCAS achieves more consistent improvements as k grows. In contrast, Vanilla-GRPO and its enhanced variants show slower growth, consistent with findings from Yue et al. (2025). The stronger performance of UCAS under the pass@k metric highlights its effectiveness, which can be attributed to differences in exploration strategies. Unlike Vanilla-GRPO, which often suffers from exploration stagnation, where the model repeatedly samples low-diversity rollouts, UCAS leverages uncertainty-aware advantage shaping to sustain diverse exploration and escape local optima. Additional pass@k results on AMC and MATH-500, reported in Appendix B.3, show the same trend on broader benchmarks.
@@ 5.1 RL for LLM Reasoning
Recent advances in reinforcement learning have transformed the training of large language models for reasoning tasks. Process reward models (PRMs) (Lightman et al., 2023; Nie et al., 2026; Wu et al., 2026) have emerged as a key innovation, providing step-level supervision that improves both efficiency and accuracy compared to outcome-only rewards. Approaches such as PRIME (Cui et al., 2025a) eliminate costly human annotation by deriving implicit process feedback, while OmegaPRM (Luo et al., 2024) leverages Monte Carlo Tree Search (MCTS) to automatically identify reasoning errors. Alongside this, DeepSeek-R1 (Guo et al., 2025) demonstrates that sophisticated reasoning can emerge purely from RL without supervised fine-tuning, enabled by GRPO, which replaces value functions with group-based baselines. These advances redefine alignment and reasoning in LLMs, positioning reinforcement learning with verifiable or process-level rewards as a scalable and principled alternative to preference-model-based RLHF.
@@ 5.2 Reinforcement Learning from Verifiable Rewards
RLVR has emerged as a scalable alternative to preference-based alignment by converting open-ended outputs into checkable signals such as mathematical correctness or unit-test pass rates (Guo et al., 2025; Yue et al., 2025; An et al., 2025; Zhou et al., 2026). While early implementations demonstrated strong gains in pass@1 accuracy, subsequent studies revealed a consistent policy entropy collapse: models rapidly concentrate probability mass on a narrow set of high-reward trajectories, diminishing output diversity and limiting exploration (Cui et al., 2025b; Hao et al., 2025). Empirical analyses show that RLVR-trained models often underperform base models on pass@ (Shao et al., 2024; Yue et al., 2025), highlighting a precision–diversity trade-off (Wu et al., 2025; Dong et al., 2025; Zhao et al., 2026).
Algorithmic responses to entropy collapse vary. Standard entropy or KL penalties provide partial remedies, though their effectiveness often depends heavily on the divergence form (Li et al., 2025). More recent uncertainty-aware approaches have sought to refine the learning signal, though with differing philosophies. For instance, SEED-GRPO (Chen et al., 2025) leverages semantic entropy to downscale updates for uncertain queries, adopting a conservative risk-mitigation strategy. In stark contrast, UCAS adopts an exploratory philosophy: we explicitly amplify rewards for correct-but-uncertain trajectories to incentivize venturing into novel reasoning domains, rather than inhibiting learning from uncertainty.
Similarly, while entropy-based shaping methods (Cheng et al., 2025) introduce indiscriminate entropy bonuses to encourage diversity, UCAS implements a conditional, two-stage mechanism. By combining a logit-space response-level confidence proxy with token-level raw logits, which we find to be a more sensitive signal of local overconfidence than post-softmax entropy, UCAS distinguishes between productive exploration and blind guessing. Unlike pure entropy-based frameworks, UCAS introduces correctness-contingent modulation, amplifying penalties for confident errors while guiding exploration through uncertainty, offering a more fine-grained solution to the entropy collapse problem than global regularization or token-level covariance control (Cui et al., 2025b).
@@ 6 Conclusion
In this work, we introduced UnCertainty-aware Advantage Shaping (UCAS), a fine-grained advantage estimation framework that leverages internal confidence signals without requiring additional reward models. By jointly modeling uncertainty at both the response and token levels, UCAS reshapes advantages to highlight critical uncertain reasoning steps, suppress overconfident errors, and preserve non-negative token-level updates on positive trajectories. Experimental results on major mathematical reasoning benchmarks show that UCAS achieves substantial performance improvements over GRPO and its enhanced variants. Analysis of the training dynamics further reveals that, as training progresses, UCAS demonstrates steadily increasing rewards, longer reasoning chains, and an entropy trajectory that first declines and then rises, reflecting stronger exploratory capability. These findings indicate that uncertainty-aware advantage shaping offers an effective pathway toward more robust reinforcement learning for large language models.
Limitations
Although UCAS demonstrates significant improvements in reasoning capabilities and exploration efficiency, we acknowledge several limitations that identify clear directions for future research. First, our experiments are exclusively focused on 1.5B and 7B parameter models, and the performance of UCAS on largerscale models has not yet been fully verified due to limitations of computing resources. Second, our method relies on self-confidence and raw logits as proxies for model uncertainty. While these internal signals are computationally efficient and effective, future work could explore alternative or complementary uncertainty metrics. Techniques such as Monte Carlo dropout, model ensembles, or semantic entropy could potentially capture different facets of model uncertainty and lead to even more refined and robust advantage shaping. Investigating these areas will be essential for understanding the broader generalizability of our approach.
Appendix
Appendix A Implementation Detials
A.1 Baselines
A.2 RL Training Configuration
We adopt the VERL framework (Sheng et al., 2024) and train our model using the optimization objective defined in Eq. 11. For both GRPO and DAPO, we use the hyperparameters in Table 4, without using entropy or KL losses. All experiments are conducted on 2 compute nodes, each equipped with 8*80GB GPUs.
| Hyperparameter | Value |
| Optimizer | AdamW |
| Actor learning rate | |
| Max prompt length | 1024 |
| Max response length | 3072 |
| Training batch size | 512 |
| Samples per prompt | 16 |
| Mini-batch size | 32 |
| Rollout temperature | 1.0 |
| Clip range , | , |
| UCAS hyperparameter , | , |